# Understanding Flow Behaviors of Supercooled Liquids by Embodying Solid-Liquid Duality at Particle Level

### v1

@@ II.2 $\tau_{\mathrm{LC}}$ in quiescent state and local response to start-up shear
The emergence of slow and fast clusters allows us to explore the heterogeneity in mechanical response. For each cluster, we calculate its shear stress by summing the atomic level stresses [69] of its constituent particles during the start-up process. Then, by averaging the results of all slow (or fast) clusters, we obtain the response of the slow (or fast) part. To mitigate the effect of thermal fluctuations, stresses are further averaged over ICE. In Fig. 5, we present the start-up shear stresses of slow part ((a) and (b)) and fast part ((c) and (d)) for both samples at different $\dot{\gamma}$. For clarity, the transient stress $\sigma(t)$ is scaled using the corresponding stable-state stress, ensuring that the resulting dimensionless stress $\sigma^{*}(t)$ is equal to $1$ at long time. We find that the stress of slow part exhibits an overshoot, a feature typically observed in glasses [70, 71]. In contrast, the stress of fast part exhibits a smooth transition to stable state without a distinct overshoot. This behavior can be described by the Maxwell model as $\sigma^{*}(t)=1-\exp(-t/\tau_{\mathrm{f}})$, where $\tau_{\mathrm{f}}$ is the Maxwell time of fast part. As $\dot{\gamma}$ increases, $\tau_{\mathrm{f}}$ is seen to decrease, meaning that the fast part becomes faster and approaches ideal viscous liquid, corresponding to the shrinkage of fast regions with increasing $\dot{\gamma}$. This observation is consistent with our previous analysis given by Figs. 3 and 4. It should be pointed out that whether there is an overshoot is not a necessary and sufficient condition to distinguish between solid-like and liquid-like responses [70, 72]. Even so, the difference between the responses of slow and fast parts is unambiguous. In the next subsection, the coexistence of solid-like and liquid-like responses will be clearly revealed in the flow under steady shear.
@@ II.3 $\tau_{\mathrm{LC}}$ in flow and local response to steady shear
For a steady shear with $\dot{\gamma}$, we calculate its $\tau_{\mathrm{LC}}$ with respect to a given time $t_{0}$ as follows: First, we switch off the shear flow at $t=t_{0}$, and measure the first-jump time from $t_{0}$ for every particle. Then, we perform the ICE average according to the configuration at $t=t_{0}$ with keeping the shear off from $t=t_{0}$. Besides $\tau_{\mathrm{LC}}$, we also calculate the first-jump time from $t=t_{0}$ in shear flow $\tau_{\mathrm{sh}}$. In calculating $\tau_{\mathrm{sh}}$, the steady shear is kept on, and the ICE average is still employed. The flow conditions for calculating $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ are illustrated in Fig. 6 (a). In the study of polymer dynamics, the abrupt cessation of flow is widely used, and its relaxation time is intimately linked to the constitutive model [6, 73]. However, for most studies of flow behaviors of particulate liquids and glasses, characteristic relaxation times are measured in the continuous flow [43, 74, 75, 76]. The times characterizing the process after the cessation of flow, though appearing in a few studies [72, 77], have not attracted enough attention. Note that $\tau_{\mathrm{sh}}$ is determined by the combined effect of the steady convection and the thermal activation from the configuration at $t=t_{0}$. In contrast, by switching off the shear, the steady convection is eliminated and, thus, $\tau_{\mathrm{LC}}$ is mainly determined by the thermal activation from the distorted configuration at $t=t_{0}$. In Fig. 6 (c), we plot the distributions of $\tau_{\mathrm{LC}}$ at equilibrium and under steady shear with various $\dot{\gamma}$ for sample A. The distribution of $\tau_{\mathrm{LC}}$ shifts to lower values as $\dot{\gamma}$ increases, corresponding to more distorted configurations at larger $\dot{\gamma}$.
@@
Figure 7 gives the $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ relations in steady shear (symbols) for both samples at various $\dot{\gamma}$ covering both the crossover regime and the shear-thinning regime. In addition, we replot the results for the start-up shear that have been shown in Fig. 3 (a) and (b) (lines). It is noteworthy that the symbols well align with the line for each $\dot{\gamma}$. This nontrivial agreement suggests that the relation between $\tau_{\mathrm{sh}}$ and $\tau_{\mathrm{LC}}$ is universal for a given $\dot{\gamma}$, irrespective of whether the flow is steady or start-up.
Figure 7: (a) and (b) show $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ obtained under steady shear (symbols) with various $\dot{\gamma}$ for sample A and sample B, respectively. In addition, the results shown in Fig. 3 (a) and (b), which are obtained by applying start-up shear, are replotted here by lines. Here, lines and symbols corresponding to the same $\dot{\gamma}$ are plotted in the same color.
To examine the likely universality of the $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ relation mentioned above, we conduct additional tests as follows. As illustrated in Fig. 6 (b), we first let the system be in a steady state with a shear rate $\dot{\gamma}_{\mathrm{i}}$. Then, we perform an abrupt change of shear rate from $\dot{\gamma}_{\mathrm{i}}$ to $\dot{\gamma}_{\mathrm{f}}$ at $t=t_{0}$. Steady shear is just a special case where $\dot{\gamma}_{\mathrm{f}}=\dot{\gamma}_{\mathrm{i}}$. Both $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ with respect to $t=t_{0}$ are calculated. In Fig. 8 (a), we show the results of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ with $\dot{\gamma}_{\mathrm{f}}=0.0011$ and several $\dot{\gamma}_{\mathrm{i}}$ ranging from $\dot{\gamma}_{\mathrm{i}}=0$ to $\dot{\gamma}_{\mathrm{i}}=0.003$ for sample A. Notice that, the conditions here include both $\dot{\gamma}_{\mathrm{i}}<\dot{\gamma}_{\mathrm{f}}$ and $\dot{\gamma}_{\mathrm{i}}\geq\dot{\gamma}_{\mathrm{f}}$. It is seen that the $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ relations at different $\dot{\gamma}_{\mathrm{i}}$ remarkably coincide for the same $\dot{\gamma}_{\mathrm{f}}$. Such coincidence is also found for different $\dot{\gamma}_{\mathrm{f}}$ (Fig. 8 (b)) and different samples (Fig. 8 (c) and (d)). The results shown in Figs. 7 and 8 suggest that the functional form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ is universal once the final shear rate $\dot{\gamma}_{\mathrm{f}}$ is given. This observation implies that the transient configuration (i.e., the configuration at $t=t_{0}$, determined by the initial shear rate $\dot{\gamma}_{\mathrm{i}}$) and the convective effect (determined by $\dot{\gamma}_{\mathrm{f}}$) independently contribute to the local relaxation.
@@
In Fig. 4, we show that the heterogeneous patterns of $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ are highly alike for the start-up shear. This likeness is also found in all other flow conditions. An example for steady shear is shown in Fig. 9. The similarity between the DH patterns of $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$, together with the universal form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ relation shown in Fig. 7 and 8, suggest an approximate one-to-one correspondence between $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$. It is interesting to note that in steady-shear cases, $\tau_{\mathrm{LC}}$ is a hidden parameter because it is not directly measurable in the continuous flow. To find $\tau_{\mathrm{LC}}$, one has to switch off the shear, which is incompatible with the condition of steady shear. On the other hand, $\tau_{\mathrm{sh}}$ can be measured directly in the flow. With the one-to-one correspondence between $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$, we can obtain the hidden $\tau_{\mathrm{LC}}$ once the distribution of $\tau_{\mathrm{sh}}$ is known.
@@
Knowing the approximate one-to-one correspondence between $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ in flow, we conjecture that the local mechanical response, which should be closely related to $\tau_{\mathrm{sh}}$, exhibits spatial heterogeneity with a pattern similar to that of $\tau_{\mathrm{LC}}$. As to the heterogeneity in response, we proposed a concept of localized elastic region (LER) [42] recently, inspired by the idea of “solidity” in supercooled liquids [78, 79, 80, 81, 82]. An LER, typically composed of hundreds of particles, is a transient, solid-like cluster capable of accumulating elastic stress in the flow of supercooled liquid. Under the steady shear, especially in the nonlinear regime, the generation, deformation and relaxation of LER are ubiquitous and successive [42, 83], contributing to the viscoelasticity of supercooled liquids. The existence of LER signifies the heterogeneity in the local response to shear: some regions behave elastically, while other regions behave as viscous liquid. Then a question arises naturally: Does LER correlate with the heterogeneous distribution of $\tau_{\mathrm{LC}}$ and the grouping of fast and slow particles? Figure 10 (a) and (b) show two slices of system A under the steady shear with $\dot{\gamma}=0.003$ at a reference time denoted as $t_{\mathrm{r}}$. The color of each particle denotes its value of $\tau_{\mathrm{LC}}$ measured at $t=t_{\mathrm{r}}$. With the method given in Ref. [83], we identify nine LERs in the system at $t=t_{\mathrm{r}}$. In Fig. 10 (a) and (b), we denote their boundaries with black dash-dot lines, and label them with L1, L2, …, L9. Meanwhile, we also plot the boundaries between the fast and slow groups determined by Eq. 2 with white solid lines. Two distinct fast clusters are found within these two slices, and are labeled with F1 and F2 in Fig. 10 (a) and (b). In Fig. 10 (c-1) – (c-8), we plot the evolution of the stresses sustained by LERs L1 – L8 (the stress of L9 is not shown due to the limited space of Fig. 10). It is seen that an LER sequentially undergoes elastic deformation, plastic yielding, and sharp disintegration that eventually transforms it into a liquid-like state [83]. An example of these three stages is denoted in Fig. 10 (c-1): Here, the beginnings of elastic deformation, plastic yielding, and disintegration are denoted as $t_{\mathrm{LER,0}}$, $t_{\mathrm{LER,1}}$, and $t_{\mathrm{LER,2}}$, respectively. In our recent work [83], we have shown that the appearance of plasticity at $t=t_{\mathrm{LER,1}}$ is accompanied by the emergence of cage-jump spots, where cage jumps highly concentrate within a few diameters of particle, inside the LER. By calculating the pseudoharmonic modes [84], we show that [83] these soft spots constitute the manifestation of the shear transformation zones (STZ) [85, 86, 87, 88]. This is very similar to amorphous solids, where STZs play as the precursor of the mechanical failure. In Fig. 10 (c-9), we show the stress evolution of the fast cluster labeled with F1. No stress accumulation is seen here. The stress is steady, similar to the macroscopic behavior of viscous liquids under steady shear.
@@
In the system, LERs are in different stages at a given time. To clarify the stages of the LERs at $t=t_{\mathrm{r}}$, we denote the reference time $t_{\mathrm{r}}$ by vertical solid lines in Fig. 10 (c-1) – (c-8). As seen in Fig. 10 (c-1) – (c-7), for LERs L1 – L7 (and L9, not shown in Fig. 10), the corresponding LER is in the elastic deformation stage (i.e., within the period of $\left[t_{\mathrm{LER,0}},t_{\mathrm{LER,1}}\right]$). In contrast, for LER L8, as seen from Fig. 10 (c-8), it is in the plastic yielding stage (i.e., within the period of $\left[t_{\mathrm{LER,1}},t_{\mathrm{LER,2}}\right]$), and is very close to the disintegration. By examining Fig. 10, we find following correlation between LER and the spatial distribution of $\tau_{\mathrm{LC}}$:
@@
(ii) For fast particles, there are two situations. One is exemplified by the fast cluster F1. It is out of all LERs. As shown in Fig. 10 (c-9), it behaves as a typical viscous liquid. The other one is exemplified by the fast cluster F2. In space, it nearly coincides with the LER L8, which is very close to its disintegration at $t=t_{\mathrm{r}}$ as shown in Fig. 10 (c-8). This coincidence is understandable, since the imminent disintegration of LER involves strong displacements of most of its constituent particles [83]. In either case, the stress of the fast cluster cannot accumulate.
Summarizing the results shown in Fig. 10, we conclude that under steady shear, the regions composed of slow particles exhibit elastic response to imposed shear, while the fast regions behave as viscous liquid. In this way, we establish a correlation between DH and local mechanical response.

### v2

@@ II.2 $\tau_{\mathrm{LC}}$ in quiescent state and local response to start-up shear
The emergence of slow and fast clusters allows us to explore the heterogeneity in mechanical response. For each cluster, we calculate its shear stress by summing the atomic level stresses [69] of its constituent particles during the start-up process. Then, by averaging the results of all slow (or fast) clusters, we obtain the response of the slow (or fast) part. To mitigate the effect of thermal fluctuations, stresses are further averaged over ICE. In figure 5, we present the start-up shear stresses of slow part ((a) and (b)) and fast part ((c) and (d)) for both samples at different $\dot{\gamma}$. For clarity, the transient stress $\sigma(t)$ is scaled using the corresponding stable-state stress, ensuring that the resulting dimensionless stress $\sigma^{*}(t)$ is equal to $1$ at long time. We find that the stress of slow part exhibits an overshoot, a feature typically observed in glasses [70, 71]. In contrast, the stress of fast part exhibits a smooth transition to stable state without a distinct overshoot. This behavior can be described by the Maxwell model as $\sigma^{*}(t)=1-\exp(-t/\tau_{\mathrm{f}})$, where $\tau_{\mathrm{f}}$ is the Maxwell time of fast part. As $\dot{\gamma}$ increases, $\tau_{\mathrm{f}}$ is seen to decrease, meaning that the fast part becomes faster and approaches ideal viscous liquid, corresponding to the shrinkage of fast regions with increasing $\dot{\gamma}$. This observation is consistent with our previous analysis given by figures 3 and 4. It should be pointed out that whether there is an overshoot is not a necessary and sufficient condition to distinguish between solid-like and liquid-like responses [70, 72]. Even so, the difference between the responses of slow and fast parts is unambiguous. In the next subsection, the coexistence of solid-like and liquid-like responses will be clearly revealed in the flow under steady shear.
@@ II.3 $\tau_{\mathrm{LC}}$ in flow and local response to steady shear
For a steady shear with $\dot{\gamma}$, we calculate its $\tau_{\mathrm{LC}}$ with respect to a given time $t_{0}$ as follows: First, we switch off the shear flow at $t=t_{0}$, and measure the first-jump time from $t_{0}$ for every particle. Then, we perform the ICE average according to the configuration at $t=t_{0}$ with keeping the shear off from $t=t_{0}$. Besides $\tau_{\mathrm{LC}}$, we also calculate the first-jump time from $t=t_{0}$ in shear flow $\tau_{\mathrm{sh}}$. In calculating $\tau_{\mathrm{sh}}$, the steady shear is kept on, and the ICE average is still employed. The flow conditions for calculating $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ are illustrated in figure 6 (a). In the study of polymer dynamics, the abrupt cessation of flow is widely used, and its relaxation time is intimately linked to the constitutive model [6, 73]. However, for most studies of flow behaviors of particulate liquids and glasses, characteristic relaxation times are measured in the continuous flow [45, 74, 75, 76]. The times characterizing the process after the cessation of flow, though appearing in a few studies [72, 77], have not attracted enough attention. Note that $\tau_{\mathrm{sh}}$ is determined by the combined effect of the steady convection and the thermal activation from the configuration at $t=t_{0}$. In contrast, by switching off the shear, the steady convection is eliminated and, thus, $\tau_{\mathrm{LC}}$ is mainly determined by the thermal activation from the distorted configuration at $t=t_{0}$. In figure 6 (c), we plot the distributions of $\tau_{\mathrm{LC}}$ at equilibrium and under steady shear with various $\dot{\gamma}$ for sample A. The distribution of $\tau_{\mathrm{LC}}$ shifts to lower values as $\dot{\gamma}$ increases, corresponding to more distorted configurations at larger $\dot{\gamma}$.
@@
Figure 7 gives the $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ relations in steady shear (symbols) for both samples at various $\dot{\gamma}$ covering both the crossover regime and the shear-thinning regime. In addition, we replot the results for the start-up shear that have been shown in figure 3 (a) and (b) (lines). It is noteworthy that the symbols well align with the line for each $\dot{\gamma}$. This nontrivial agreement suggests that the relation between $\tau_{\mathrm{sh}}$ and $\tau_{\mathrm{LC}}$ is universal for a given $\dot{\gamma}$, irrespective of whether the flow is steady or start-up.
Figure 7: (a) and (b) show $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ obtained under steady shear (symbols) with various $\dot{\gamma}$ for sample A and sample B, respectively. In addition, the results shown in figure 3 (a) and (b), which are obtained by applying start-up shear, are replotted here by lines. Here, lines and symbols corresponding to the same $\dot{\gamma}$ are plotted in the same color.
To examine the likely universality of the $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ relation mentioned above, we conduct additional tests as follows. As illustrated in figure 6 (b), we first let the system be in a steady state with a shear rate $\dot{\gamma}_{\mathrm{i}}$. Then, we perform an abrupt change of shear rate from $\dot{\gamma}_{\mathrm{i}}$ to $\dot{\gamma}_{\mathrm{f}}$ at $t=t_{0}$. Steady shear is just a special case where $\dot{\gamma}_{\mathrm{f}}=\dot{\gamma}_{\mathrm{i}}$. Both $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ with respect to $t=t_{0}$ are calculated. In figure 8 (a), we show the results of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ with $\dot{\gamma}_{\mathrm{f}}=0.0011$ and several $\dot{\gamma}_{\mathrm{i}}$ ranging from $\dot{\gamma}_{\mathrm{i}}=0$ to $\dot{\gamma}_{\mathrm{i}}=0.003$ for sample A. Notice that, the conditions here include both $\dot{\gamma}_{\mathrm{i}}<\dot{\gamma}_{\mathrm{f}}$ and $\dot{\gamma}_{\mathrm{i}}\geq\dot{\gamma}_{\mathrm{f}}$. It is seen that the $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ relations at different $\dot{\gamma}_{\mathrm{i}}$ remarkably coincide for the same $\dot{\gamma}_{\mathrm{f}}$. Such coincidence is also found for different $\dot{\gamma}_{\mathrm{f}}$ (figure 8 (b)) and different samples (figure 8 (c) and (d)). The results shown in figures 7 and 8 suggest that the functional form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ is universal once the final shear rate $\dot{\gamma}_{\mathrm{f}}$ is given. This observation implies that the transient configuration (i.e., the configuration at $t=t_{0}$, determined by the initial shear rate $\dot{\gamma}_{\mathrm{i}}$) and the convective effect (determined by $\dot{\gamma}_{\mathrm{f}}$) independently contribute to the local relaxation.
@@
In figure 4, we show that the heterogeneous patterns of $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ are highly alike for the start-up shear. This likeness is also found in all other flow conditions. An example for steady shear is shown in figure 9. The similarity between the DH patterns of $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$, together with the universal form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ relation shown in figures 7 and 8, suggest an approximate one-to-one correspondence between $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$. It is interesting to note that in steady-shear cases, $\tau_{\mathrm{LC}}$ is a hidden parameter because it is not directly measurable in the continuous flow. To find $\tau_{\mathrm{LC}}$, one has to switch off the shear, which is incompatible with the condition of steady shear. On the other hand, $\tau_{\mathrm{sh}}$ can be measured directly in the flow. With the one-to-one correspondence between $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$, we can obtain the hidden $\tau_{\mathrm{LC}}$ once the distribution of $\tau_{\mathrm{sh}}$ is known.
@@
Knowing the approximate one-to-one correspondence between $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ in flow, we conjecture that the local mechanical response, which should be closely related to $\tau_{\mathrm{sh}}$, exhibits spatial heterogeneity with a pattern similar to that of $\tau_{\mathrm{LC}}$. As to the heterogeneity in response, we proposed a concept of localized elastic region (LER) [44] recently, inspired by the idea of “solidity” in supercooled liquids [78, 79, 80, 81, 82]. An LER, typically composed of hundreds of particles, is a transient, solid-like cluster capable of accumulating elastic stress in the flow of supercooled liquid. Under the steady shear, especially in the nonlinear regime, the generation, deformation and relaxation of LER are ubiquitous and successive [44, 83], contributing to the viscoelasticity of supercooled liquids. The existence of LER signifies the heterogeneity in the local response to shear: some regions behave elastically, while other regions behave as viscous liquid. Then a question arises naturally: Does LER correlate with the heterogeneous distribution of $\tau_{\mathrm{LC}}$ and the grouping of fast and slow particles? Figure 10 (a) and (b) show two slices of system A under the steady shear with $\dot{\gamma}=0.003$ at a reference time denoted as $t_{\mathrm{r}}$. The color of each particle denotes its value of $\tau_{\mathrm{LC}}$ measured at $t=t_{\mathrm{r}}$. With the method given in Ref. [83], we identify nine LERs in the system at $t=t_{\mathrm{r}}$. In figure 10 (a) and (b), we denote their boundaries with black dash-dot lines, and label them with L1, L2, …, L9. Meanwhile, we also plot the boundaries between the fast and slow groups determined by equation 2 with white solid lines. Two distinct fast clusters are found within these two slices, and are labeled with F1 and F2 in figure 10 (a) and (b). In figure 10 (c-1) – (c-8), we plot the evolution of the stresses sustained by LERs L1 – L8 (the stress of L9 is not shown due to the limited space of figure 10). It is seen that an LER sequentially undergoes elastic deformation, plastic yielding, and sharp disintegration that eventually transforms it into a liquid-like state [83]. An example of these three stages is denoted in figure 10 (c-1): Here, the beginnings of elastic deformation, plastic yielding, and disintegration are denoted as $t_{\mathrm{LER,0}}$, $t_{\mathrm{LER,1}}$, and $t_{\mathrm{LER,2}}$, respectively. In our recent work [83], we have shown that the appearance of plasticity at $t=t_{\mathrm{LER,1}}$ is accompanied by the emergence of cage-jump spots, where cage jumps highly concentrate within a few diameters of particle, inside the LER. By calculating the pseudoharmonic modes [84], we show that [83] these soft spots constitute the manifestation of the shear transformation zones (STZ) [85, 86, 87, 88]. This is very similar to amorphous solids, where STZs play as the precursor of the mechanical failure. In figure 10 (c-9), we show the stress evolution of the fast cluster labeled with F1. No stress accumulation is seen here. The stress is steady, similar to the macroscopic behavior of viscous liquids under steady shear.
@@
In the system, LERs are in different stages at a given time. To clarify the stages of the LERs at $t=t_{\mathrm{r}}$, we denote the reference time $t_{\mathrm{r}}$ by vertical solid lines in figure 10 (c-1) – (c-8). As seen in figure 10 (c-1) – (c-7), for LERs L1 – L7 (and L9, not shown in figure 10), the corresponding LER is in the elastic deformation stage (i.e., within the period of $\left[t_{\mathrm{LER,0}},t_{\mathrm{LER,1}}\right]$). In contrast, for LER L8, as seen from figure 10 (c-8), it is in the plastic yielding stage (i.e., within the period of $\left[t_{\mathrm{LER,1}},t_{\mathrm{LER,2}}\right]$), and is very close to the disintegration. By examining figure 10, we find following correlation between LER and the spatial distribution of $\tau_{\mathrm{LC}}$:
@@
(ii) For fast particles, there are two situations. One is exemplified by the fast cluster F1. It is out of all LERs. As shown in figure 10 (c-9), it behaves as a typical viscous liquid. The other one is exemplified by the fast cluster F2. In space, it nearly coincides with the LER L8, which is very close to its disintegration at $t=t_{\mathrm{r}}$ as shown in figure 10 (c-8). This coincidence is understandable, since the imminent disintegration of LER involves strong displacements of most of its constituent particles [83]. In either case, the stress of the fast cluster cannot accumulate.
Summarizing the results shown in figure 10, we conclude that under steady shear, the regions composed of slow particles exhibit elastic response to imposed shear, while the fast regions behave as viscous liquid. In this way, we establish a correlation between DH and local mechanical response.